OpenStudy (anonymous):

How do you prove cos(2x) = ... using the unit circle?

OpenStudy (dumbcow):

the trig identity ? cos(2x) = cos^2(x) - sin^2(x)

OpenStudy (anonymous):

Right, how is it derived using the unit circle?

OpenStudy (dumbcow):

hmm i know it comes from angle sum formula cos(a+b) = cos(a)cos(b) - sin(a)sin(b) ...don't remember how they derive it though.. some fancy geometry i believe

OpenStudy (anonymous):

right, it is the fancy geometry I am looking for